Vapour pressure of water at 298 is 17.51mm, lowering of vapour pressure of sugar solution is 0.614mm.calculate RLVP
 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
the vapour pressure of solvent p1o = 17.51 mm Hg
 
lowering of vapour pressure = 0.0614 mm Hg
 
hence RLVP = 0.0614/ 17.51  = 3.51 *10^-3